# Dynamic Diffusion Schrödinger Bridge in Astrophysical Observational Inversions

### v2

@@ Abstract
We study Diffusion Schrödinger Bridge (DSB) models in the context of dynamical astrophysical systems, specifically tackling observational inverse prediction tasks within Giant Molecular Clouds (GMCs) for star formation. We introduce the Astro-DSB model, a variant of DSB with the pairwise domain assumption tailored for astrophysical dynamics. By investigating its learning process and prediction performance in both physically simulated data and in real observations (the Taurus B213 data), we present two main takeaways. First, from the astrophysical perspective, our proposed paired DSB method improves interpretability, learning efficiency, and prediction performance over conventional astrostatistical and other machine learning methods. Second, from the generative modeling perspective, probabilistic generative modeling reveals improvements over discriminative pixel-to-pixel modeling in Out-Of-Distribution (OOD) testing cases of physical simulations with unseen initial conditions and different dominant physical processes. Our study expands research into diffusion models beyond the traditional visual synthesis application and provides evidence of the models’ learning abilities beyond pure data statistics, paving a path for future physics-aware generative models which can align dynamics between machine learning and real (astro)physical systems. Code available at https://github.com/L-YeZhu/AstroDSB.
@@ 1 Introduction
Diffusion models (DMs) (Sohl-Dickstein et al., 2015; Ho et al., 2020; Song et al., 2020), sometimes also known as score-based generation models (SGMs), have become a popular probabilistic modeling approach in machine learning for many data generation applications (Dhariwal and Nichol, 2021; Ho et al., 2022; Rombach et al., 2022). A recent line of work improves the vanilla design by reformulating the mapping trajectory learning in the context of the Schrödinger bridge problem (Schrödinger, 1932), which allows for imposing domain-specific restrictions on the learned trajectory, leading to Diffusion Schrödinger Bridge models (DSB) (Wang et al., 2021; De Bortoli et al., 2021; Shi et al., 2023). These DSB models are theoretically elegant and have proven effective in vision-related applications such as image generation and translation (Liu et al., 2023; Gushchin et al., 2024). However, we argue that the true power of this reformulation lies in the fact that it allows for diffusion models to capture the underlying data dynamics in scientific domains where the state transitions themselves hold physical meaning.
We consider the problem setting of astrophysical observational inverse prediction, specifically in the context of Giant Molecular Cloud (GMCs) formation and evolution leading to star formation — most stars in our Galaxy form from such systems (Fig. 1). In this setting we are able to observe snapshots, either from simulations or from real observations, but these only represent the boundary states of complex physical processes. Effectively modeling these processes requires not only being able to predict the outcomes under familiar conditions (in-distribution) but also being able to adapt to changes in both the initial physical conditions and/or in the dominant dynamics (out-of-distribution). This is in stark contrast to the relative simplicity of the vision tasks like image transition, which involve direct mappings between discrete data points (e.g., pixel configurations) with no underlying temporal or physical evolution. Astrophysics offers an exciting application because it explores some of the most complex, dynamic, and high-dimensional systems in the universe, from the turbulent collapse of interstellar clouds to the birth of stars and the evolution of galaxies. These phenomena generate diverse observational data, and accurately modeling these hidden processes not only tests the limits of machine learning but also enhances our understanding of the universe.
In this work, we introduce and study Astro-DSB, a Diffusion Schrödinger Bridge model tailored to the astrophysics domain by aligning the diffusion model states with the dynamics governing GMC evolution. We first evaluate the quality of this approach on density and magnetic field prediction compared to the ground truth from simulation data. Then, most importantly, we examine its behavior under distribution shifts using both physically simulated data and real astrophysical observations. In contrast to conventional vision tasks, where OOD is often heuristically defined, astrophysical systems provide scientifically meaningful definitions of distributional shifts through variations in initial conditions and dominant physical processes (Wang and Abel, 2009; Wu et al., 2015). These configurations reflect realistic physical diversity, making them especially suitable for testing distribution-level generalization. Through this evaluation, we uncover the practical boundaries of probabilistic modeling and its advantage in domains where the underlying dynamics, rather than just data statistics, govern the observations.
@@ 2.2 Inverse astrostatistical predictions for star formation
Traditionally in astronomy, such observational inverse prediction problems have been commonly approached via astrostatistical methods, which often combine theoretical simplifications with statistical priors (detailed in Sec. 4). For instance, molecular density is commonly estimated using two- or three-component power-law conversion derived under idealized physical assumptions about emission properties and cloud structure (Bisbas et al., 2021, 2023). While effective within limited regimes, these methods tend to lack flexibility and generalization ability when applied to systems governed by varying physical conditions, making their performance unreliable in unseen testing cases.
@@ 3.1 Problem statement and preliminary in astronomy and ML
Given a physical initial condition $\mathbf{x}_{0}$, we assume that the GMCs evolve under a set of governing physical laws as modeled in previous astrophysical studies (Wu et al., 2015, 2020; Hsu et al., 2023). This results in an equilibrium physical state $\mathbf{x}_{1}$, which we denote as $\mathbf{x}_{1}=\mathcal{H}(\mathbf{x}_{0})$, where $\mathcal{H}$ is a highly complex, nonlinear function incorporating self-gravity, magnetic fields, radiative heating, and cooling processes. While the forward process $\mathcal{H}$ is typically well-defined in controlled simulations, direct access to $\mathbf{x}_{1}$ is rarely available in practice. Instead, we often only observe a limited projection $\mathbf{y}$ of the true physical state. Therefore, our objective is to infer the full distribution $\mathbf{x}_{1}$ from partial observable $\mathbf{y}$, featuring a class of observational inverse prediction problems that lie at the core of many tasks in astrostatistical modeling.
We assume an observational model in the form of:
@@
where $\mathcal{F}$ represents the observational process (e.g., line-of-sight integration, instrumental response) and $\varepsilon$ denotes corresponding observational noise in the measurement.
Given this formulation, our core task becomes a conditional probabilistic inference problem, where the goal is to estimate the posterior distribution $p(\mathbf{x}_{1}|\mathbf{y})$.
@@
We are given two distributions $p_{0},p_{1}$ and a reference path measure $\mathbb{Q}\in\mathcal{P}(\mathcal{C})$, with $\mathcal{C}=\mathrm{C}([0,1],\mathbb{R}^{d})$ defined as the space of continuous functions from time space $[0,1]$ to $\mathbb{R}^{d}$. The dynamic SB problem then seeks to find a path measure $\mathbb{P}^{SB}\in\mathcal{P}(\mathcal{C})$ such that:
@@
SB can be formalized as an entropy-regularized Optimal Transport (OT) problem describing the following forward and backward SDEs :
@@ 3.2 Diffusion Schrödinger bridge for astrophysical observational prediction (Astro-DSB)
Figure 2: Illustration pipeline of our proposed Astro-DSB method. In the training stage, we propose to learn the DSB model under the pairwise matching assumption with observational noises alignment $\varepsilon$ and enhancement observables $\mathbf{y}$. In inference, we adopt an aggregated scalable sampling strategy to infer physical distributions from larger observables.
@@
[…]
@@
Following conventional image translation (Liu et al., 2023) and other physical tasks such as field reconstruction (Li et al., 2025) that take a clean data input $\mathbf{y}$, we note that explicitly modeling the noise perturbation term $\epsilon$ in observables $\mathbf{y}$ in Eq. 5 offers a simple yet critical alignment to physical assumptions in observational prediction scenarios described in Eq. 1. In this work, we assume that the noise introduced from the observation process follows a weighted Gaussian, denoted as $\epsilon\sim\lambda\mathcal{N}(0,I)$. Following statistical conventions in astrophysical studies, we set $\lambda=0.1$ throughout the experiments. This modification ensures that our Astro-DSB framework respects the uncertainty structure of constrained observables, refining the targeted conditional posterior distribution to be $p(\mathbf{x}_{1}|\mathbf{y}-\varepsilon)$.
@@
We enhance the learning process by incorporating the clean observable input $\mathbf{y}$ as an additional intermediate condition at each time step $t$. This modification to the vanilla I2SB framework facilitates the learning of the score function $\epsilon_{\theta,t}$ implemented through the U-Net architecture (Ronneberger et al., 2015) by consistently “reminding” the model of the observational constraint throughout the trajectory. Empirically, we find this design is critical for achieving robust prediction performance within the DSB framework and helps obtain both lower loss values and a faster convergence speed.
@@
$\mathcal{F}_{\theta,t}=||\epsilon_{\theta,t}(\mathbf{x}_{t},\mathbf{y},t;\theta)-\frac{\mathbf{y}-\varepsilon-\mathbf{x}_{1}}{\sigma_{t}}||,$ (6)
@@
For inference, in contrast to the padding-based techniques used in prior works (Li et al., 2025), we design an aggregated sampling strategy to restore higher-resolution prediction results. Given a large testing observable $\mathbf{y}_{obs}$ (e.g., real-world survey data), we partition $\mathbf{y}_{obs}$ into overlapping patches $\mathbf{y}^{i}_{obs}$ that match the training resolution. We then independently run inference on each patch to obtain corresponding predictions $\tilde{\mathbf{x}}^{i}_{1}$, and aggregate across all patches to reconstruct the final predicted distribution $\tilde{\mathbf{x}}_{1}$. Fig. 2 illustrates the described pipeline.
@@ 4.1 Astrophysical data
Similar to previous work (Xu et al., 2023b, 2025), our training data are drawn from synthetic magnetohydrodynamics (MHD) simulations. Specifically, we follow the simulation setup from Wu et al. (2020) and Hsu et al. (2023), which is conducted using the MUSCL-Dedner method and HLLD Riemann solver in the adaptive mesh refinement (AMR) code Enzo (Dedner et al., 2002; Wang and Abel, 2009; Bryan et al., 2014). These simulations model key physical processes including self-gravity, magnetic fields, and radiative heating and cooling, based on a photodissociation model (Wu et al., 2015). We consider the molecular density ($\rho$, measured in the number of H nuclei $n_{H}/cm^{-3}$) and magnetic field strength ($\mathcal{B}$, measured in $\mu G$) as the primary quantities for prediction in this work. For molecular density, we collected data from 7179 simulation trials, split into 5707 training and 1472 testing cases following an 8:2 ratio. For magnetic field strength, we obtained 19100 training and 6370 testing cases. All data are pre-processed by a logarithmic transformation followed by normalization to [0,1], and resized to $128\times 128$ patches.
@@
Additionally, we generate 1680 independent tests with different dominant physical process assumptions as the OOD testing cases to evaluate generalization. Specifically, unlike the ID MHD simulations code ENZO (Bryan et al., 2014) initialized with colliding and non-colliding molecular clouds, OOD data are produced using a different MHD code, ORION2 (Li et al., 2021), which excludes gravity but includes driven turbulence, and employs a different set of initial conditions. While the fundamental MHD equations are similar between two simulation codes, the physical processes and initial setups differ significantly, thus clearly distinguishing the ID and OOD cases under physically meaning shifts.
Synthetic and real observables.
@@
After generating the simulation data, we define the observables $\mathbf{y}$ following conventions in astrophysical studies. For molecular density, we adopt the projected mass surface density maps (also known as the “column density", measured in $N_{H}/cm^{-2}$) as the observables $\mathbf{y}_{\rho}$. For the magnetic field, the synthetic observable $\mathbf{y}_{\mathcal{B}}$ consists of the column density, the orientation angles of the dust continuum polarization vector, and line-of-sight (LOS) nonthermal velocity dispersion. In addition to these synthetic observables, we also evaluate the proposed Astro-DSB on real observational data from the Taurus B213 region, using column density maps provided by  Palmeirim et al. (2013). B213 is one of closest star-forming filamentary structures located in the Taurus molecular cloud (Li and Goldsmith, 2012). Unlike the synthetic data where ground-truth distributions are available, the Taurus B213 observation $\mathbf{y}_{B213}$ lacks a precise reference for physical states, and is instead constrained by indirect measurements from complementary observations. This real-data evaluation demonstrates the applicability of Astro-DSB in practical scenarios where ground truth is unavailable, and highlights its potential for scientific discovery in observational settings.
4.2 Prior methods and evaluation setup
Astrostatistical methods.

### v3

@@ Abstract
We study Diffusion Schrödinger Bridge (DSB) models in the context of dynamical astrophysical systems, specifically tackling observational inverse prediction tasks within Giant Molecular Clouds (GMCs) for star formation. We introduce the Astro-DSB model, a variant of DSB with the pairwise domain assumption tailored for astrophysical dynamics. By investigating its learning process and prediction performance in both physically simulated data and in real observations (the Taurus B213 data), we present two main takeaways. First, from the astrophysical perspective, our proposed paired DSB method improves interpretability, learning efficiency, and prediction performance over conventional astrostatistical and other machine learning methods. Second, from the generative modeling perspective, probabilistic generative modeling reveals improvements over discriminative pixel-to-pixel modeling in Out-Of-Distribution (OOD) testing cases of physical simulations with unseen initial conditions and different dominant physical processes. Our study expands research into diffusion models beyond the traditional visual synthesis application and provides evidence of the models’ learning abilities beyond pure data statistics, paving a path for future physics-aware generative models which can align dynamics between machine learning and real (astro)physical systems.  [^1] Code available at https://github.com/L-YeZhu/AstroDSB
@@ 1 Introduction
Diffusion models (DMs) (Sohl-Dickstein et al., 2015; Ho et al., 2020; Song et al., 2020), sometimes also known as score-based generation models (SGMs), have become a popular probabilistic modeling approach in machine learning for many data generation applications (Dhariwal and Nichol, 2021; Ho et al., 2022; Rombach et al., 2022). A recent line of work improves the vanilla design by reformulating the mapping trajectory learning in the context of the Schrödinger bridge problem (Schrödinger, 1932), which allows for imposing domain-specific restrictions on the learned trajectory, leading to Diffusion Schrödinger Bridge models (DSB) (Wang et al., 2021; De Bortoli et al., 2021; Shi et al., 2023). These DSB models are theoretically grounded and have proven effective in vision-related applications such as image generation and translation (Liu et al., 2023; Gushchin et al., 2024). However, their modeling capabilities and limits could extend beyond vision, offering a new opportunity to test these generative formulations in the context of real-world physical systems, where data are governed by complex dynamics and observational constraints rather than pixel statistics.
In this work, we study a real-world physical problem, focusing on a key class of inverse prediction tasks in astrophysics, particularly within the context of star formation (McKee and Ostriker, 2007; Larson, 2003; Madau and Dickinson, 2014). Specifically, we aim to better understand and predict the physical states within Giant Molecular Clouds (GMCs) (Shu et al., 1987), which is a major component of the interstellar medium (ISM) (Spitzer Jr, 2008) that directly regulates star formation rates. The goal of our task is to infer true physical distributions, such as density or magnetic fields, from partial observables, as illustrated in Fig. 1.
As a critical and active research area in astrophysics, multiple methods have been developed over the past decades to tackle this challenge. Given that observational inverse prediction problems are inherently intractable, early approaches often relied on astrostatistical methods such as power-law fitting (Maschberger and Kroupa, 2009), which combine theoretical simplifications with observational statistical priors (detailed in Sec. 2.2 and Sec. 4.2). However, these methods typically suffer from large prediction errors and limited generalization on out-of-distribution (OOD) testing cases under shifted physical initial conditions, due to often oversimplified physical assumptions and restricted capacity for modeling complex non-linearities.
@@
Given the increasing capability of neural networks in modeling large-scale data with highly complicated nonlinearities, the astrophysical community has recently started to explore machine learning approaches (Yang et al., 2024; Xu et al., 2020a) to overcome previous limitations. Among multiple ML architectures, conditional diffusion generative models (DGMs) (Sohl-Dickstein et al., 2015; Ho et al., 2020; Song et al., 2020) stand out as recent state-of-the-art (SOTA) frameworks with large improvements in prediction accuracy over conventional methods for several critical physical quantities, such as GMCs density distribution (Xu et al., 2023b) and magnetic field strengths (Xu et al., 2025). These methods typically take the observables as input conditions and output the predicted physical distribution maps through the distribution mapping ability of DGMs trained on simulation data. However, despite the enhanced prediction accuracy, such conditional diffusion models share a few constraints, including a physics-ML misalignment due to the Gaussian prior from the original DM formulation (also detailed in Sec. 3.2), as well as the known expensive computational cost for learning DGMs.
To this end, we introduce Astro-DSB, a variant of Diffusion Schrödinger Bridge model tailored to astrophysics with better physics-ML state alignment and improved learning efficiency. Specifically, compared to existing DGMs-based methods that learn a mapping between the Gaussian prior and the true physical distribution with the observable as the condition, Astro-DSB removes the reliance on the Gaussian prior from the DGMs formulation and learn the direct mapping between observables and ground-truth physical states. This design also helps to accelerate the learning efficiency by 75% compared to the previous DGM-based SOTA. More concretely, Astro-DSB leverages the intrinsic deterministic mapping property between the observation and ground truth physical state and adopts a paired assumption between two boundary distributions  [^2] Terminology note: The term “paired DSB” is used here in the ML sense, which refers to the sample-level conditioning between observables and physical states, rather than in the classical unpaired Schrödinger Bridge formulation based on probabilistic coupling.. Within this high-level Schrödinger bridge ML framework, we further integrate the observables as an enhanced condition to facilitate the bridge learning process. Intuitively, this can be mathematically considered as learning a conditional Doob’s h-transform, thus differentiating from the vanilla paired DSB models in computer vision tasks (Liu et al., 2023) and scientific applications such as protein designs (Didi et al., 2023).
Our comprehensive experiments on both molecular density and magnetic field tasks validate the effectiveness of our proposed method. In addition to the comparable or superior SOTA performance in ID testing cases, we also examine the performance under distribution shifts using both physically simulated data and real astrophysical observations from Taurus B213. In contrast to conventional vision tasks, where OOD is often heuristically defined, astrophysical systems provide scientifically meaningful definitions of distributional shifts through variations in initial conditions and dominant physical processes (Wang and Abel, 2009; Wu et al., 2015). These configurations reflect realistic physical diversity, making them especially suitable for testing distribution-level generalization. Through this evaluation, we uncover the practical boundaries of probabilistic modeling and its advantages in scientific domains such as astrophysics.
@@ 2.2 Inverse astrostatistical predictions for star formation
Traditionally in astronomy, such observational inverse prediction problems have been commonly approached via astrostatistical methods, which often combine theoretical simplifications with statistical priors (also in Sec. 4). For instance, molecular density is commonly estimated using two- or three-component power-law conversion derived under idealized physical assumptions about emission properties and cloud structure (Bisbas et al., 2021, 2023). While effective within limited regimes, these methods tend to lack generalization ability when applied to systems governed by varying physical conditions and dynamics, making their performance unreliable in unseen testing cases.
@@ 3.1 Problem statement and preliminary in astronomy and ML
Given a physical initial condition $\mathbf{x}_{0}$, we assume that the GMCs evolve under a set of governing physical laws as modeled in previous astrophysical studies (Wu et al., 2015, 2020; Hsu et al., 2023). This results in an equilibrium physical state $\mathbf{x}_{1}$, which we denote as $\mathbf{x}_{1}=\mathcal{H}(x_{0})$, where $\mathcal{H}$ is a highly complex, nonlinear function incorporating self-gravity, magnetic fields, radiative heating, and cooling processes. While the forward process $\mathcal{H}$ is typically well-defined in controlled simulations, direct access to $\mathbf{x}_{1}$ is rarely available in practice. Instead, we often only observe a limited projection $\mathbf{y}$ of the true physical state. Therefore, our objective is to infer the full distribution $\mathbf{x}_{1}$ from partial observable $\mathbf{y}$, featuring a class of observational inverse prediction problems that lie at the core of many tasks in astrostatistical modeling. We further assume an observational model in the form of:
@@
where $\mathcal{F}$ represents the observational process (e.g., line-of-sight integration, instrumental response) and $\varepsilon$ denotes corresponding observational noise in the measurement. Given this formulation, our core task becomes a conditional probabilistic inference problem, where the goal is to estimate the posterior distribution $p(\mathbf{x}_{1}|\mathbf{y})$.
@@
Consider two distributions $p_{0},p_{1}$ and a reference path measure $\mathbb{Q}\in\mathcal{P}(\mathcal{C})$, the dynamic SB problem seeks to find a path measure $\mathbb{P}^{SB}\in\mathcal{P}(\mathcal{C})$ such that:
@@
with $\mathcal{C}=\mathrm{C}([0,1],\mathbb{R}^{d})$ defined as the space of continuous functions from time space $[0,1]$ to $\mathbb{R}^{d}$. SB can be formalized as an entropy-regularized OT problem describing the following forward and backward SDEs:
@@ 3.2 Diffusion Schrödinger bridge for astrophysical observational prediction (Astro-DSB)
Figure 2: Illustration pipeline of our proposed Astro-DSB method. In the training stage, we propose to learn the DSB model under the pairwise matching assumption with observational noises alignment $\varepsilon$ and enhancement observables $\mathbf{y}$. In inference, we adopt an aggregated scalable sampling strategy to infer physical distributions from larger observables, such as Taurus B213.
@@
where $\sigma^{2}_{t}:=\int^{t}_{0}\beta_{\tau}d\tau$ and $\bar{\sigma}^{2}_{t}:=\int_{t}^{1}\beta_{\tau}d\tau$ denote accumulated variances. This forms the base design of our proposed Astro-DSB. However, we observe that this baseline framework alone is insufficient for accurately predicting true physical distributions from constrained observational inputs.
@@
Compared to conventional image translation (Liu et al., 2023) and other physical tasks such as field reconstruction (Li et al., 2025) that take a clean data input $\mathbf{y}$, we note that explicitly modeling the noise perturbation term $\epsilon$ in observables $\mathbf{y}$ in Eq. 5 offers a simple yet critical alignment to physical assumptions in observational prediction scenarios described in Eq. 1. In this work, we assume that the noise introduced from the observation process follows a weighted Gaussian, denoted as $\epsilon\sim\lambda\mathcal{N}(0,I)$. Following statistical conventions in astrophysical studies, we set $\lambda=0.1$ throughout the experiments. This modification ensures that our Astro-DSB framework respects the uncertainty structure of constrained observables, refining the targeted conditional posterior distribution to be $p(\mathbf{x}_{1}|\mathbf{y}-\varepsilon)$.
@@
Compared to the vanilla I2SB framework, we further propose to enhance the learning process by incorporating the clean observable input $\mathbf{y}$ as an additional intermediate condition at each time step $t$. This modification aims to facilitate the learning of the score function $\epsilon_{\theta,t}$, implemented through U-Net architecture (Ronneberger et al., 2015), by consistently “reminding” the model of the observational constraint throughout the trajectory. Empirically, we find this design critical for achieving robust prediction performance within the DSB framework and helps simultaneously obtain lower loss values with a faster convergence speed.
@@
[…]
@@
For inference, in contrast to the padding-based techniques used in prior works (Li et al., 2025), we design an aggregated sampling strategy to restore higher-resolution prediction results. Given a large testing observable $\mathbf{y}_{obs}$ (e.g., real-world survey data), we partition $\mathbf{y}_{obs}$ into overlapping patches $\mathbf{y}^{i}_{obs}$ that match the training resolution. We then independently run inference on each patch to obtain corresponding predictions $\tilde{\mathbf{x}}^{i}_{1}$, and aggregate across all patches to reconstruct the final predicted distribution $\tilde{\mathbf{x}}_{1}$. Fig. 2 illustrates the described pipeline, with detailed algorithms in the appendices.
@@ 4.1 Astrophysical data
Magnetohydrodynamics simulations. Similar to previous work (Xu et al., 2023b, 2025), our training data are drawn from synthetic magnetohydrodynamics (MHD) simulations. Specifically, we follow the simulation setup from Wu et al. (2020) and Hsu et al. (2023), which are conducted using the MUSCL-Dedner method and HLLD Riemann solver in the adaptive mesh refinement (AMR) code Enzo (Dedner et al., 2002; Wang and Abel, 2009; Bryan et al., 2014). These simulations model key physical processes including self-gravity, magnetic fields, radiative heating, and cooling, based on a photodissociation model (Wu et al., 2015). We consider the molecular density ($\rho$, measured in the number of H nuclei $n_{H}/cm^{-3}$) and magnetic field strength ($\mathcal{B}$, measured in $\mu G$) as the primary quantities for prediction in this work. For molecular density, we collected data from 7179 simulation trials, split into 5707 training and 1472 testing cases following an 8:2 ratio. For magnetic field strength, we obtained 19100 training and 6370 testing cases. All data are pre-processed by a logarithmic transformation followed by normalization to [0,1], and resized to $128\times 128$ patches.
@@
Additionally, we generate 1680 independent tests with different dominate physical process assumptions as the OOD testing cases to evaluate generalization. Specifically, unlike the ID MHD simulations code ENZO (Bryan et al., 2014) initialized with colliding and non-colliding molecular clouds, OOD data are produced using a different MHD code, ORION2 (Li et al., 2021), which excludes gravity but includes driven turbulence, and employs a different set of initial conditions. While the fundamental MHD equations are similar between two simulation codes, the physical processes and initial setups differ significantly, thus clearly distinguishing the ID and OOD cases under physically meaning shifts.
@@
Synthetic and real observables. After generating the simulation data, we define the observables $\mathbf{y}$ following conventions in astrophysical studies. For molecular density, we adopt the projected mass surface density maps (also known as the “column density", measured in $N_{H}/cm^{-2}$) as the observables $\mathbf{y}_{\rho}$. For the magnetic field, the synthetic observable $\mathbf{y}_{\mathcal{B}}$ consists of the column density, the orientation angles of the dust continuum polarization vector, and line-of-sight (LOS) nonthermal velocity dispersion. In addition to these synthetic observables, we also evaluate the proposed Astro-DSB on real observational data from the Taurus B213 region, using column density maps provided by  Palmeirim et al. (2013). B213 is one of closest star-forming filamentary structures located in the Taurus molecular cloud (Li and Goldsmith, 2012). Unlike the synthetic data where ground-truth distributions are available, the Taurus B213 observation $\mathbf{y}_{B213}$ lacks a precise reference for physical states, and is instead constrained by indirect measurements from complementary observations. This real-data evaluation demonstrates the applicability of Astro-DSB in practical scenarios where ground truth is unavailable, and highlights its potential for scientific discovery in observational settings.
4.2 Comparing methods and evaluations